6x + 5y = 51
7x + 3y = 51

Answer by mananth(16946)

6x + 5y = 51
7x + 3y = 51
multiply equation 1 by -7 and equation 2 by 6
we get
-42x-35y=-357
42x+18y=306
add the two equations
-35y+18y-42x+42x=-351+306
-17y=-51
/-17
y= -51/-17
y= 51/17
y=3
...
plug the value of y in any equation
6x+5*3=51
6x+15=51
6x=51-15
6x = 36
x= 6